A particle is thrown with a velocity of u m/s. It passes A and B at time t1 = 1s and t2 = 3s. The value of u is (g=10 m/s2).

Grade:

2 Answers

Vaani
6 years ago
Final time = 4 sec 4= 2u sino/h4 = 2u Sin30/10 U= 40(Total time = 3+1=4)(Angle of elevation = y/x)
Ex
5 years ago
At max. height, t=2s (as this position is midway between t=1s and t=3s)
t=usin@/g( to reach the max. height)
t= usin30°/g
2= u*1/2*10
U=2*2*10
U=40ms–¹.